Can you please help me solve this equation. I have tried numerous time, getting frustrated.
Write the equation of the line which passes through the point (2,1) and is parallel to the line with equation y=5/2x+6

Answer by longjonsilver(2297)

very simple topic really.

The critical thing you need to remember is y=mx+c. Any straight line will/can be written like this, where m is the gradient and c is the y-intercept.

So, we have the line y = (5/2)x + 6... it is written in standard form, so its gradient is 5/2.

The line we are asked for, is parallel to this one. Parallel means the same gradient, so our line has m=5/2, so we have...

y = (5/2)x + c

Now we need to know c. Well to do that, we need to know x and y, ie a point on the line...which surprise surprise, we do, namely (2,1). So, putting these into the equation, we get

1 = (5/2)(2) + c
--> 1 = 5 + c
--> -4 = c

so, that is it, our equation is y = (5/2)x - 4, which we may write without fractions as 2y = 5x - 8.

jon.